OpenStudy (fefetasprites):

A new photocopier can make 72 copies in 2 min. When an older photocopier is operating, the two photocopiers together can make 72 copies in 1.5 min. How long would it take the older photocopier to make 72 copies alone? I'm not sure how to set this up, can someone please help?

OpenStudy (anonymous):

I set it up like this it's very blurry and I apologize for that! I did new= 72/2 and New + Old= 72/1.5 I finished the problem so if you need me to walk it through it, I can!

OpenStudy (fefetasprites):

Thank you for you help!